What is the expected value with perfect information of the following decision table?

States of nature

Alternative

S1

S2

Probability

0.4

0.6

Option 1

10,000

30,000

Option 2

5,000

45,000

Option 3

-4,000

60,000

A) 5,000         B) 10,000         C) 40,000         D) 60,000         E) 70,000
Answer #1

With perfect information, the best outcome of an alternative will be selected.

In S1, option 1 is the highest (at 10,000) and so it will be selected.

In S2, option 3 is the highest (at 60,000) and so it will be selected.

Expected value = probability*value

= (0.4*10,000) + (0.6*60,000)

= 4,000+36,000 = 40,000

Answer is option C.